Circular No. 8319

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 F4 (BRADFIELD) On Mar. 28, A. C. Beresford, Myrtle Bank, South Australia, reported

the visual discovery of a possible new comet close to the horizon in twilight by William A. Bradfield, during his search for sungrazing comets, from observations made with a 0.25-m reflector on Mar. 23.43 and 24.42 UT; further attempts to observe it had failed until the last few days. The delay in reporting was due to difficulties in matching the sketch of observations with star charts. Below are additional observations that have been reported following requests to several observers.

2004 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 23.43  1^{h}50.^{m}0  −1^{o}26′  8  Bradfield
24.42  1 51.0  −1 25  _{”}
Apr. 8.40  2 08.5  −1 36  _{”}
9.368 2 09.1  −1 24  5  McNaught
11.3461 2 08 32.^{s}74 −0 56 37′′.8  4  Lovejoy
12.3427 2 07 39.50 −0 30 59.2  3.3  _{”}

W. A. Bradfield (Yankalilla, S.A.). Original positions given for equinox

1950.0. Magnitude guessed. R. H. McNaught and T. M. Smith (Siding Spring Observatory). Strongly

condensed blue coma with diameter ∼ 0′.5; hint of a tail pointing up from the horizon. Comet low in evening twilight; 20×120 binoculars. Positional uncertainty stated as ± 5′. In-focus magnitude comparable to two nearby stars of mag 5.6. Also observed visually on Apr. 12.358, again almost stellar, blue, and with a short (perhaps 2′-long?) tail pointing up to the left at 10 degrees from vertical.

T. Lovejoy (Thornlands, Qld., Australia). λ = 153^{o}.267, ϕ = −27^{o}.55, h = 30 m. On Apr. 11, nineteen co-added CCD images with Canon 300D camera + 100-mm f/4 camera lens; fairly obvious tail 0^{o}.5 long.

The following very preliminary parabolic orbital elements are from the above observations.

T = 2004 Apr. 17.12 TT  ω = 332.^{o}49
Ω = 222.66 }2000.0
q = 0.1690 AU  i = 63.21
2004 April 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green